Inessential Menu Ordering

1 Where to find the menus

Menus are stored in the following sections: MIT - On-Campus section has on-campus dining options and convenience stores. Deli & Burgers section has delis, sandwich shops, burgers, grill joints. Fish & Fowl section has seafood and poultry places. Italian & Pizza section has Italian food and pizza parlors. Mexican section has Mexican, Tex-Mex, and Southwestern-themed places. Chinese - Popular section has extra copies of up-to-date menus of popular Chinese restaurants

only. These may be removed from the folder, but return them. The other Chinese folder has one copy of each of these menus.

Chinese - Other section has Cantonese, Hunan, Hong Kong, Mandarin, and Szechuan cuisine.

Popular Chinese restaurants should have a single copy of their menus in this folder. Indian section has Indian and Middle Eastern restaurants. Other Asian section has Japanese, Korean, Thai, Vietnamese etc. restaurants. Miscellaneous mostly has diners, desert places, and other places that go across categories. Old Menus section has menus that are more than four years old, but might be worth keeping

anyway. Menus are alphabetical by name in each section.

2 How to treat the menus

• Keep the binder, folders, and menus neat.

• Do not remove any blue folders from this binder. If you do, you will need to replace the folder and its menus so other people can easily find menus. Don’t be lazy: fix your mistake immediately.

• Do not remove any menus bound in this binder when browsing. In general, there is only one copy and we do not want it lost, spilled on, mutilated, etc.

• Do pass around copies of Popular Chinese menus for browsing and order taking. Keep them clean and put them back neatly when you are done.

• If you visit or order from a restaurant that has an old menu, request a new menu so we can keep prices up to date. Request multiple copies of menus of popular restaurants. Place them in an appropriate section of the binder, in alphabetical order, so they are easy to find.

• Put the date (month and year) in the upper-left corner of the first page of any menu you add to the binder. Remove older or undated menus.

• Store useful coupons in the pocket inside the binder. Discard expired coupons if you notice them.

This document can be found in /afs/sipb/admin/text/labels/menus.tex and was last updated on March 20, 1998.